Let A and B be two events with P(A)=0.5, P(B)=0.4, P(A ∩ B')=0.2. Find P(B'/A U B).
1
Expert's answer
2020-06-22T18:14:29-0400

To find :


P(B'/ AUB)


Formula: conditional probability,P(A/B)= P(A and B)/ P(B)


Solution : Using conditional probability formula ,


P(B'/AUB)=P(B' and AUB)/P(AUB)


P(B' and AUB)= P(An B')= 0.2


P(An B)= P(A)-P(AnB')=0.5-0.2=0.3


P(AUB)= P(A)+P(B)=P(A n B)= 0.5+0.4-0.3=0.6


Hence, P(B'/AUB)=0.2/0.6=0.33


attached Venn diagram for reference.